OpenStudy (anonymous):

Two urns each contain green balls and red balls. Urn I contains 10 green balls and 8 red balls, and Urn II contains 3 green balls and 10 red balls. A ball is drawn from each urn. Find P(red and red).

OpenStudy (anonymous):

\[\frac{ 8}{18} * \frac{ 3}{13}\\\\ \frac{ 4}{9} * \frac{ 3}{13}\\\\ \frac{ 4}{39} \]

OpenStudy (anonymous):

Great, thank you!